OpenStudy (moongazer):

How do you get the quotient of this. (x^6+y^6 )÷(x+y)

OpenStudy (lgbasallote):

i dont think it's factorable :/ you'll have to go with long division sorry

OpenStudy (anonymous):

Ha ha ha.. After a long struggle you wrote this??

OpenStudy (anonymous):

You can get..

OpenStudy (lgbasallote):

hmm it's complicated with long division too

OpenStudy (moongazer):

What should I do?

OpenStudy (anonymous):

See, \[(x^2)^3 + (y^2)^3 = (x^2+y^2)(x^4 + y^4 - x^2y^2)\]

OpenStudy (moongazer):

I got that one. That is from the sum of two cubes. but what should I do next?

OpenStudy (anonymous):

I am not finding any strong way to do that.. But let me go by long division method..

OpenStudy (moongazer):

did you got something?

OpenStudy (anonymous):

\[\large \color{green}{Quotient = x^5 -x^4y + x^3y^2-x^2y^3 + xy^4 - y^5}\] And, \[\large \color{blue}{Remainder = 2y^6}\]

OpenStudy (anonymous):

I found this by using Long Division Method.. I think it could be more simpler to solve this if it was: \(x^6 - y^6\)

OpenStudy (moongazer):

How did you divide it?

OpenStudy (anonymous):

It is a long division method.. How can I show you I am not getting..

OpenStudy (moongazer):

could you please show me step by step on how did you get the quotient. please... :)

OpenStudy (moongazer):

@waterineyes

OpenStudy (anonymous):

I show you step by step but It may take time..

OpenStudy (moongazer):

Thank you :)

OpenStudy (moongazer):

I am really confused with this :)

OpenStudy (anonymous):